OpenStudy (anonymous):

Can someone help me? I have to write the equation of the line using the point-slope formula (0, 4) and (2, 3). 0 – 2 4 – 3 Change the equation to slope-intercept form. Can someone explain how to do this?

OpenStudy (anonymous):

Calculate slope by (y2-y1)/(x2-x1) = -1/2

OpenStudy (anonymous):

slope = (3-4)/(2-0) intercept => at point x=0 you see that y is 4. This means you now can create the formula: y = slope*x + intercept y = -1/2 x + 4

OpenStudy (anonymous):

Thank you.